Consider a US firm that imports components from China, and adds value by assembling these components into a finished product in the US. Assume that the US firm always pays in RMB (Chinese Remnibi) for the components. Suppose the cost of imported components is RMB 25, while domestic assembly costs to finish the product are $5, and the current exchange rate is RMB5/$.

1)

Case A: If the US firm was selling 100% of the finished product output in the domestic US

market at $12/unit, what is the margin of the US firm as a % of revenue?

2) Case B: If the US firm was exporting 100% of the finished product output to Eurozone markets

and gets paid €12/unit in Euros, and the exchange rate is $1/Euro

what is the margin of the US firm as % of revenue?

3) Case C: In case A above, if the exchange rate changed to RMB 4.5/$, and all the other numbers

and variables remain constant, what is the margin of the US firm as % of revenue?

4)

Case D:

In case B above, if the exchange rate changed to $1.1/€, and all the other numbers and variables remain constant

(RMB5/$), what is the new margin of the US firm as % of revenue
Answer #1

$1 is equal to 5 RMB hence $5 is equal to 25 RMB total cost production is equal to 5 dollar + 5 Dollar equal to $10

As good sold at $12 margin is 2$

Hence % = 2/12*100=16.67%

B) $1 is equal to 1 Euro hence there will be no impact on the margin 16.67% that means same as above in part a

C) as now exchange rate $1 is equal to 4.5 RMB

Goods are produced at 25 RMB this implies they are produced that 5.55 additional cost of 5 and total 10.55

Hence margin 12-10.55=1.45

Margin = 1.45/12*100= 12.083%

D) as our product is sold at 12€ which is 13.2 $ (12*1.1)

Total cost still 5$+5$=10$, hence margin =13.2-10=3.2

%=3.2/13.2*100= 24.24%